# Towards Distillation-Resistant Large Language Models: An Information-Theoretic Perspective

### v1

@@ 4.2 Transformation Matrix Optimization
Next, we analyze the proposed matrix training objective through the lens of the Information Bottleneck (IB) principle to resolve the optimization problem. In particular, the following property reveals that minimizing CMI actually induces an IB objective.
@@
The conditional mutual information $\mathcal{I}(X;Z|Y)$ is a special case of Information Bottleneck objective with $\beta=1$ in Equation 3, i.e.,
@@
  • Minimizing $\mathcal{I}(X;Z)$. This compresses redundant contextual information in the teacher logits $Z$ that originates from the input $X$, weakening the dependency between inputs and outputs. As a consequence, $Z$ becomes less informative about $X$, which enhances the resistance against KD algorithms.
  • Maximizing $\mathcal{I}(Z;Y)$. This preserves task-relevant information and maintains the model’s utility. Notably, it encourages $Z$ to concentrate toward its class centroid (i.e., increasing the correct label’s logit while suppressing others), which also eliminates contextual information and brings defense benefits.
Training Loss Design. Motivated by the two sub-tasks, we design two complementary loss functions to achieve the CMI minimization. We begin with analyzing the second term $\mathcal{I}(Z;Y)$, which can be expanded into
@@
Note that $\mathcal{H}(Y)$ is constant, thus maximizing $\mathcal{I}(Z;Y)$ is equivalent to minimizing the conditional entropy $\mathcal{H}(Y|Z)$, which measures the residual uncertainty about the ground-truth label $Y$ after observing logits $Z$. Although directly computing $\mathcal{H}(Y|Z)$ is intractable, we can derive a consistent upper bound as follows.
@@
For maximizing $\mathcal{I}(X;Z)$, previous studies on image classification typically compute its value (Ye et al., 2024) or estimate it via variational upper bounds (IB变分). However, directly estimating $\mathcal{I}(X;Z)$ for autoregressive LLMs is fundamentally infeasible, since both $X$ and $Z$ are in high-dimensional spaces, and their underlying distributions are difficult to approximate. To handle this challenge, we note that in KD scenarios, the information of $X$ residing in $Z$ is ultimately exploited through how $Z$ instructs the student’s output, which in practice materializes in the form of gradients that update the student’s parameters during training. Hence, the learning signals conveyed by the gradients serve as the effective carrier of the information of $X$ residing in $Z$, which motivates the following proposition:
@@
$\bm{g}$ $=\nabla_{\theta}\frac{1}{l}\sum_{t=1}^{l}\Big[-(1-\alpha)\cdot\log\sigma\big(f_{\textnormal{S}}(q,o_{<t})\big)_{o_{t}}$ (8)
$+\alpha\cdot\mathbb{KL}\Big(\sigma\big(f_{\textnormal{T}}(q,o_{<t})\big)\Big\|\sigma\big(f_{\textnormal{S}}(q,o_{<t})\big)\Big)\Big].$
@@
With the above loss terms, the training objective for learning the transformation matrix $\mathcal{M}$ is formulated as:
$\mathcal{L}_{\mathcal{M}}=\mathcal{L}_{\text{CE}}+\lambda\cdot\mathcal{L}_{\text{grad}},$ (9)
$=-\frac{1}{l}\sum_{t=1}^{l}\log\sigma\big(f_{\textnormal{T}}(q,o_{<t})\big)_{o_{t}}+\lambda\cdot\frac{\langle\bm{g},\bm{g}^{\prime}\rangle}{\|\bm{g}\|_{F}\|\bm{g}^{\prime}\|_{F}},$
where $\lambda>0$ is the scaling factor of $\mathcal{L}_{\mathrm{grad}}$, $\langle\cdot,\cdot\rangle$ is the dot operation, and $\|\cdot\|_{F}$ denotes the Frobenius norm.
Transformation Matrix Optimization. The pseudocode of training algorithm is presented in Appendix  C. Notably, a direct implementation of $\mathcal{M}$ as a dense $|\mathcal{V}|\times|\mathcal{V}|$ matrix is computationally prohibitive due to the large vocabulary size. To mitigate this, we parameterize $\mathcal{M}$ via a low-rank decomposition $\mathcal{M}=\mathbf{E}+\mathbf{AB}$, where $\mathbf{E}$ denotes the identity matrix, and $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times r}$ and $\mathbf{B}\in\mathbb{R}^{r\times|\mathcal{V}|}$ are learnable matrices with decomposition rank $r$. Moreover, directly computing full gradients renders the optimization of $\mathcal{L}_{\text{grad}}$ intractable due to their high dimensionality. To this end, we adopt the gradients of the student model’s second-to-last layer as a surrogate in the loss calculation, as they contain rich task-relevant information and closely align with the model’s predicting behaviors. Inspired by LoRA, we initialize $\mathbf{A}$ from $\mathcal{N}(0,1)$ and initialize $\mathbf{B}=\mathbf{0}$, so that $\mathcal{M}$ starts from $\mathbf{E}$ and ensures $Z^{\prime}=Z$ at the beginning of training. This initialization yields more stable gradients and improves optimization effectiveness (Hu et al., 2022).
@@ 5.1 Experimental Setup
Models and Datasets. Note that logit-based KD requires teachers and students to share the same tokenization to ensure one-to-one correspondence in token semantics. Hence, we select models from the same series to avoid engineering issues from vocabulary mismatches and better focus on defensive analysis. We consider two widely used LLM families, Qwen2.5 (Yang et al., 2024) and Llama 3 (Grattafiori et al., 2024). For the Qwen family, we use Qwen2.5-7B as the teacher model and Qwen2.5-3B-Instruct as the surrogate student. Distillation is performed on multiple student models with different scales and capacities, including Qwen2.5-3B/1.5B/0.5B. For the Llama family, we use Llama-3.1-8B as the teacher model and Llama-3.2-3B-Instruct as the surrogate student. The defensive performance is then evaluated on student models Llama-3.2-3B and Llama-3.2-1B.
Following prior work (Savani et al., 2025; Li et al., 2025), we consider three widely used benchmarks, GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2020), and MATH (Hendrycks et al., 2021). GSM8K is a math problem dataset with annotated reasoning chains; we use the standard split with 7,473 training and 1,319 evaluation examples. MMLU is a large-scale benchmark for evaluating world knowledge and problem-solving. We randomly sample 15,000 pairs as training data and evaluate on a test set of 14,042 examples. The MATH is a more challenging math benchmark, which contains 7,500 training examples with step-by-step solutions. Following (Xia et al., 2025), we evaluate performance on the MATH-500 (Lightman et al., 2023) dataset.
Distillation Baselines. We evaluate our method against four logit-based KD baselines that have exhibited competitive performance across a wide range of tasks: the classical logit-based KD (Hinton et al., 2015) with forward KLD, reverse KLD (RKLD) (Gu et al., 2024), AlphaNet (Wang et al., 2021), and the state-of-the-art ABKD (Wang et al., 2025).
Table 1: Performance of the teacher model under the defense.
Teacher Model GSM8K MMLU
vanilla Ours vanilla Ours
Qwen2.5-7B 80.89 79.83 73.85 72.84
Llama-3.1-8B 55.95 54.44 65.20 63.99
Figure 3: Visualization comparison of defensive and original teacher outputs using Qwen2.5-7B on GSM8K
Table 2: Distillation Performance of different student models under different distillation approaches (%).
@@
Implementation Details. We train the proposed transformation matrix for 5 epochs at a batch size of 32. For GSM8K and MATH, the learning rate is set to $5\cdot 10^{-7}$, while for MMLU is $5\cdot 10^{-6}$. The matrix decomposition rank is set to 32. The learning rate follows a linear warm-up schedule over the first 10% of total training steps and then decays to zero using cosine annealing. We optimize the matrix using AdamW with a weight decay of 0.01 and default momentum coefficients $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We adopt the implementation and corresponding hyperparameter settings provided in (Wang et al., 2025), using $\alpha=0.1$ and $\beta=0.9$ for AlphaNet, and $\alpha=0.1$ and $\beta=0.8$ for ABKD. In Sec. 5.4, we explore the influence of the loss factor $\lambda$.
For a stable and reproducible evaluation, we conduct all experiments using greedy decoding. More details including student distillation configuration are in Appendix B.2.
5.2 Performance Evaluation
We first report the accuracy of defensive teacher models in Table 1. By maximizing the mutual information between $Z^{\prime}$ and $Y$ using $\mathcal{L}_{\mathrm{CE}}$, the learned matrix effectively preserves teacher accuracy. Figure 3 provides outputs generated from the original and the transformed teacher. As observed, our defensive teacher produces fluent and coherent responses while preserving task correctness. This contrasts with existing sequence-level distillation defenses that disrupt sampling distributions and induce texts with degraded quality.
@@ 5.2 Performance Evaluation
We also observe that our defense has a stronger effect on GSM8K than MMLU, which we attribute to GSM8K’s longer multi-step mathematical reasoning, where perturbations accumulate along the reasoning trajectory and exert a more pronounced effect. By contrast, MMLU mainly involves multi-choice questions that LLMs typically resolve with a direct option selection, making it less sensitive to such disruptions. Due to space limits, we provide results on surrogate models in Appendix D and visualization of student outputs under distillation in Appendix E.
@@
Figure 5: Accuracy of teacher and distilled student models under varying values of the matrix loss parameter $\lambda$ and decomposition rank $r$.
Loss Landscape. To reveal the underlying mechanism of our approach, we then analyze the training dynamics of the overall loss $\mathcal{L}_{\mathcal{M}}$ and the gradient deviation loss $\mathcal{L}_{\mathrm{grad}}$. As shown in Figure 5, the cosine similarity term $\mathcal{L}_{\text{grad}}$ starts around $1$, indicating that the gradients of transformed logits are initialized to be identical to original ones (i.e., $\mathcal{M}=\mathbf{E}$ initially). In the first 300 steps, the overall loss $\mathcal{L}_{\mathcal{M}}$ decreases while $\mathcal{L}_{\text{grad}}$ remains nearly unchanged, implying that this stage mainly optimizes $\mathcal{L}_{\mathrm{CE}}$. Afterwards, $\mathcal{L}_{\text{grad}}$ rapidly drops to negative values, indicating that the learned matrix induces a large gradient angular deviation.
5.3 Distillation on the MATH benchmark
Apart from the above two benchmarks, we further evaluate on the more challenging MATH benchmark, which contains significantly more difficult math problems. Since the Llama series performs poorly on this dataset, making the evaluation less informative, we conduct experiments using the stronger Qwen series.
Table 3: Performance of Qwen2.5 student models on MATH-500 under different distillation algorithms (%).
@@ 5.3 Distillation on the MATH benchmark
After applying the learned matrix, the teacher’s accuracy decreases only marginally from 53.8% to 50.4%. In contrast, Table 3 shows that distilled students suffer a far more pronounced degradation. E.g., under KD with Qwen2.5-3B, the accuracy drops by a notable value of 6.8%. This again confirms the effectiveness and universality of our CMI-inspired defense across different datasets.
5.4 Ablation Study
We conduct ablation analysis on Qwen using GSM8K under standard KD. More ablations are in Appendix D.
Ablation on the matrix loss parameter $\lambda$. $\lambda$ is the scaling factor that balances the contributions of $\mathcal{L}_{\mathrm{CE}}$ and $\mathcal{L}_{\text{grad}}$ during matrix training in Eq. (LABEL:eq:final-objective). The left subfigure of Figure 5 shows that when increasing $\lambda$, there is a general trend of decreasing student accuracy, which confirms that $\mathcal{L}_{\text{grad}}$ further reduces distillation-relevant information. However, overly large $\lambda$ also leads to a noticeable drop in the teacher’s accuracy. We then choose $\lambda=1$ to achieve a better trade-off between defense strength and utility preservation.
Ablation on the decomposition rank $r$ of $\mathcal{M}$. The right subfigure of Figure 5 reveals a trade-off between teacher utility and defensive efficacy. When the rank $r$ is small, the limited capacity mainly prioritizes minimizing $\mathcal{L}_{\mathrm{CE}}$, preserving teacher accuracy but insufficient defensive effects. In contrast, larger $r$ allows more capacity for optimizing $\mathcal{L}_{\mathrm{grad}}$, enhancing defense while potentially compromising teacher accuracy. To balance defense effectiveness and model utility, we adopt $r=32$ as the default setting.
@@ 6 Conclusion
In this paper, we investigate the challenging problem of defending LLMs against unauthorized logit-based KD. We characterize the guidance signals in the teacher logits via conditional mutual information, based on which we propose a post-processing approach that learns a transformation matrix using two theory-inspired loss functions. By refining the output logits, the proposed method degrades the performance of distilled student models while preserving teacher utility, effectively protecting the intellectual property. We highlight that, unlike prior defenses against sequence-level distillation, our method makes a pioneering step in defending against tricky logit-based KD techniques. We hope this work advances the community’s understanding of LLM distillation and motivates more robust distillation defenses.
Impact Statement
This paper presents one of the earliest systematic attempts to protect the intellectual property of large language models against logit-based knowledge distillation, a particularly challenging problem that has been largely overlooked by prior work. Experimental results demonstrate that the proposed method is effective across a wide range of settings while preserving the utility of the teacher model. Moreover, this work aims to deepen the community’s understanding of the mechanisms underlying knowledge distillation. Overall, we believe this research advances machine learning by improving the privacy and security of LLM. There are many potential societal consequences of this line of research, most of which are related to the secure deployment, responsible model sharing, and intellectual property protection of LLMs, and none that we believe require specific discussion here.

### v2

@@ 4.2 Transformation Matrix Optimization
Next, we analyze the proposed matrix training objective through the lens of the Information Bottleneck (IB) principle to resolve the optimization problem. In particular, the following property reveals that minimizing $\mathcal{I}(X;Z|Y)$ in this scenario actually induces an IB objective.
@@
The conditional mutual information $\mathcal{I}(X;Z|Y)$ is a special case of Information Bottleneck objective with $\beta=1$ in eq.˜3, i.e.,
@@
  • Minimizing $\mathcal{I}(X;Z)$. This compresses redundant contextual information in teacher logits $Z$ that originates from the input $X$, weakening the dependency between inputs and outputs. As a result, $Z$ becomes less informative about $X$, which enhances the resistance against KD algorithms.
  • Maximizing $\mathcal{I}(Z;Y)$. This preserves task-relevant information to maintain model utility. Notably, it encourages $Z$ to concentrate toward its class centroid (i.e., increasing the correct label’s logit while suppressing others), which also reduces contextual information and brings defense benefits.
Training loss design. Motivated by the two sub-tasks, we design two complementary loss functions to minimize CMI. We begin with analyzing the second term $\mathcal{I}(Z;Y)$, which can be expanded into
@@
Note that $\mathcal{H}(Y)$ is constant for a fixed dataset, thus maximizing $\mathcal{I}(Z;Y)$ is equivalent to minimizing the conditional entropy $\mathcal{H}(Y|Z)$, which measures the residual uncertainty about the ground-truth label $Y$ after observing logits $Z$. Although directly computing $\mathcal{H}(Y|Z)$ is generally intractable, we can derive a theoretically grounded and tractable upper bound as follows.
@@
For minimizing $\mathcal{I}(X;Z)$, previous studies on image classification typically compute its value Ye et al. (2024) or estimate it via variational upper bounds IB变分. However, directly estimating $\mathcal{I}(X;Z)$ for autoregressive LLMs is fundamentally infeasible, since both $X$ and $Z$ are in high-dimensional spaces, and their underlying distributions are difficult to approximate. To handle this challenge, we note that in KD scenarios, the information of $X$ residing in $Z$ is ultimately exploited through how $Z$ instructs the student’s output, which in practice materializes in the form of gradients that update the student’s parameters during training. Hence, the learning signals conveyed by the gradients serve as the effective carrier of the information of $X$ residing in $Z$, which motivates the following proposition:
@@
$\bm{g}=\nabla_{\theta}\frac{1}{l}\sum_{t=1}^{l}\Big[-(1-\alpha)\cdot\log\sigma\big(f_{\textnormal{S}}(q,o_{<t})\big)_{o_{t}}+\alpha\cdot\mathbb{KL}\Big(\sigma\big(f_{\textnormal{T}}(q,o_{<t})\big)\Big\|\sigma\big(f_{\textnormal{S}}(q,o_{<t})\big)\Big)\Big].$ (10)
@@
With these loss terms, the optimization objective for the transformation matrix $\mathcal{M}$ is formulated as:
$\mathcal{L}_{\mathcal{M}}=\mathcal{L}_{\text{CE}}+\lambda\cdot\mathcal{L}_{\text{grad}}=-\frac{1}{l}\sum_{t=1}^{l}\log\sigma\big(f_{\textnormal{T}}(q,o_{<t})\big)_{o_{t}}+\lambda\cdot\frac{\langle\bm{g},\bm{g}^{\prime}\rangle}{\|\bm{g}\|_{F}\|\bm{g}^{\prime}\|_{F}},$ (11)
where $\lambda>0$ is the scaling factor, $\langle\cdot,\cdot\rangle$ is the dot operation, and $\|\cdot\|_{F}$ denotes the Frobenius norm.
Transformation matrix optimization. We provide the pseudocode of the complete training algorithm in Appendix  D. Notably, a direct implementation of $\mathcal{M}$ as a dense $|\mathcal{V}|\times|\mathcal{V}|$ matrix is computationally prohibitive due to the large vocabulary size. To mitigate this issue, we adopt a parameter-efficient formulation by decomposing $\mathcal{M}$ into a low-rank structure, $\mathcal{M}=\mathbf{E}+\mathbf{AB}$, where $\mathbf{E}$ denotes the identity matrix, and $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times r}$ and $\mathbf{B}\in\mathbb{R}^{r\times|\mathcal{V}|}$ are learnable matrices with decomposition rank $r$. Moreover, directly computing full gradients renders the optimization of $\mathcal{L}_{\text{grad}}$ intractable due to their high dimensionality. To this end, we adopt the gradients of the student model’s second-to-last layer as a surrogate in the loss calculation, as they contain rich task-relevant information and closely align with the model’s predicting behaviors. Inspired by LoRA, we initialize $\mathbf{A}$ from $\mathcal{N}(0,1)$ and initialize $\mathbf{B}=\mathbf{0}$, so that $\mathcal{M}$ starts from $\mathbf{E}$ and ensures $Z^{\prime}=Z$ at the beginning of training. This initialization yields more stable gradients and improves optimization effectiveness Hu et al. (2022).
@@ 5.1 Experimental Setup
Models and datasets. Note that logit-based KD requires teachers and students to share the same tokenization to ensure one-to-one correspondence in token semantics. Hence, we select models from the same series to avoid engineering issues from vocabulary mismatches and better focus on defensive analysis. Following Savani et al. (2025); Li et al. (2025), we use two widely used LLM families, Qwen2.5 Yang et al. (2024) and Llama 3 Grattafiori et al. (2024). For the Qwen family, we use Qwen2.5-7B as the teacher model and Qwen2.5-3B-Instruct as the surrogate student. Distillation is performed on multiple student models with different scales and capacities, including Qwen2.5-3B/1.5B/0.5B. For the Llama family, we use Llama-3.1-8B as the teacher model and Llama-3.2-3B-Instruct as the surrogate student. The defensive performance is then evaluated on student models Llama-3.2-3B and Llama-3.2-1B.
Following prior work Savani et al. (2025); Li et al. (2025), we consider three widely used benchmarks, GSM8K Cobbe et al. (2021), MMLU Hendrycks et al. (2020), and MATH Hendrycks et al. (2021). GSM8K is a math problem dataset with annotated reasoning chains; we use the standard split with 7,473 training and 1,319 evaluation examples. MMLU is a large-scale benchmark for evaluating world knowledge and problem-solving. We randomly sample 15,000 pairs as training data and evaluate on a test set of 14,042 examples. The MATH is a more challenging math benchmark, which contains 7,500 training examples with step-by-step solutions. Following Xia et al. (2025), we evaluate performance on the MATH-500 Lightman et al. (2023) dataset.
Distillation baselines. We evaluate our method against four logit-based KD baselines that have exhibited competitive performance across a wide range of tasks: the classical logit-based KD Hinton et al. (2015) with forward KLD, reverse KLD (RKLD) Gu et al. (2024), AlphaNet Wang et al. (2021), and the state-of-the-art ABKD Wang et al. (2025).
Implementation details. We train the transformation matrix for 5 epochs at a batch size of 32. For GSM8K and MATH, the learning rate (lr) is set to $5\cdot 10^{-7}$, while for MMLU is $5\cdot 10^{-6}$. The matrix decomposition rank is set to 32. The lr follows a linear warm-up schedule over the first 10% of total training steps and then decays to zero using cosine annealing. We optimize the matrix using AdamW with a weight decay of 0.01 and default momentum coefficients $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We adopt the implementation and corresponding hyperparameter settings provided in Wang et al. (2025), using $\alpha=0.1$ and $\beta=0.9$ for AlphaNet, and $\alpha=0.1$ and $\beta=0.8$ for ABKD. In Sec. 5.3, we explore the influence of the loss factor $\lambda$. For a stable and reproducible evaluation, we conduct all experiments using greedy decoding. More details including student distillation configuration are in Appendix B.2.
5.2 Performance Evaluation
Table 1: Distillation performance of different student models under different distillation approaches.
@@ 5.2 Performance Evaluation
We first report the accuracy of defensive teacher models in Figure 3(a). By maximizing mutual information between $Z^{\prime}$ and $Y$ using $\mathcal{L}_{\mathrm{CE}}$, the learned matrix effectively preserves teacher accuracy. Figure 3(b) shows visualizations of generated outputs. As observed, our defensive teacher produces fluent and coherent responses while preserving task correctness. This contrasts with existing sequence-level distillation defenses that disrupt sampling distributions and induce texts with degraded quality.
Figure 3: Performance of the teacher model under the defense and visualization comparison of defensive and original teacher outputs using Qwen2.5-7B on GSM8K.
@@
We also observe that our defense has a stronger effect on GSM8K than MMLU, which we attribute to GSM8K’s longer multi-step mathematical reasoning, where perturbations accumulate along the reasoning trajectory and exert a more pronounced effect. By contrast, MMLU mainly involves multi-choice questions that LLMs typically resolve with a direct option selection, making it less sensitive to such disruptions. Due to space limits, we provide results on surrogate models in Appendix C and visualization of student outputs under distillation in Appendix F.
Figure 4: Accuracy of teacher and distilled student models on the GSM8K dataset under varying values of the matrix loss parameter $\lambda$ and decomposition rank $r$.
@@
Table 2: Performance of Qwen2.5 student models on MATH-500 under different distillation methods.
@@
Distillation on the MATH benchmark. Beyond the above two benchmarks, we further evaluate our method on the more challenging MATH dataset, which comprises substantially more difficult mathematical problems. As the Llama series performs poorly on this dataset, which makes the evaluation less informative, we instead conduct experiments using the stronger Qwen series to obtain more meaningful results.
After being equipped with the learned matrix, the teacher’s accuracy exhibits only a marginal decrease, dropping from 53.8% to 50.4%. In contrast, as shown in Table 2, the distilled students suffer a far more pronounced degradation. E.g., under KD with Qwen2.5-3B, the accuracy drops by a notable value of 6.8%. This again confirms the effectiveness and universality of our CMI-inspired defense strategy across different datasets.
5.3 Ablation Study and Loss Landscape Analysis
This part first provides ablation analysis on Qwen using GSM8K under standard KD. We then present a loss landscape analysis to investigate the optimization process. More ablations are in Appendix C.
Ablation on the matrix loss parameter $\lambda$. $\lambda$ is an important scaling factor that balances the contributions of $\mathcal{L}_{\mathrm{CE}}$ and $\mathcal{L}_{\text{grad}}$ during matrix training in Eq. (11). The left subfigure of Figure 5 shows that when increasing $\lambda$, there is a general trend of decreasing student accuracy, which confirms that $\mathcal{L}_{\text{grad}}$ further suppresses distillation-relevant information and strengthens the defense. However, overly large $\lambda$ also leads to a noticeable drop in the teacher’s accuracy. We then choose $\lambda=1$ to achieve a better trade-off between defense strength and utility preservation.
Ablation on the decomposition rank $r$ of $\mathcal{M}$. The right subfigure of Figure 5 reveals a trade-off between teacher utility and defensive efficacy. When the rank $r$ is small, the limited model capacity primarily focuses on minimizing $\mathcal{L}_{\mathrm{CE}}$, thereby preserving the teacher’s accuracy but yielding insufficient defensive effects against knowledge extraction. In contrast, larger $r$ allows more capacity for optimizing $\mathcal{L}_{\mathrm{grad}}$, enhancing defense while potentially compromising teacher accuracy. Therefore, we adopt $r=32$ as the default setting to balance defense effectiveness and model utility.
Loss landscape. To better understand the underlying mechanism of our approach, we analyze the training dynamics of the overall loss $\mathcal{L}_{\mathcal{M}}$ and the gradient deviation loss $\mathcal{L}_{\mathrm{grad}}$. As shown in Figure 5, the cosine similarity term $\mathcal{L}_{\text{grad}}$ starts around $1$, indicating that the gradients of transformed logits are initialized to be identical to original ones (i.e., $\mathcal{M}=\mathbf{E}$ initially). During the first 300 training steps, the overall loss $\mathcal{L}_{\mathcal{M}}$ steadily decreases while $\mathcal{L}_{\text{grad}}$ remains nearly unchanged, implying that this phase is dominated by the optimization of $\mathcal{L}_{\mathrm{CE}}$. As training progresses, $\mathcal{L}_{\text{grad}}$ rapidly drops to negative values, indicating that the learned matrix induces a significant deviation in gradient angular.
@@ 6 Conclusion
In this paper, we investigate the challenging problem of defending LLMs against unauthorized logit-based KD from an information-theoretical perspective. We characterize the guidance signals in the teacher logits via conditional mutual information, based on which we propose a post-processing approach that learns a transformation matrix using two theory-inspired loss functions. By refining the output logits, the proposed method degrades the performance of distilled student models while preserving teacher utility, effectively protecting the intellectual property. We highlight that, unlike prior defenses against sequence-level distillation, our method makes a pioneering step in defending against tricky logit-based KD techniques. We hope this work advances the community’s understanding of LLM distillation and motivates more robust distillation defenses.